OpenStudy (anonymous):

PLEASE HELP Which of the following systems of equations has infinitely many solutions? 3x + 3y = 11; 6y = 6x - 22 8x - 7 = 6y; 12y - 14 = 4x 7y = x + 3; -3x + 21y = 9 5x + 7y = 4; 14y = 10x + 8

OpenStudy (anonymous):

anything parrallel has infinite solutions. Parallel and lay on top of eachother.

OpenStudy (anonymous):

so iuld be the first one?

OpenStudy (neo92):

nope. to be parallel their gradients must be the same. arrange the eqn in the form \[ax+by+c=0\] then, \[gradient(m)=-a/b\] now try for the answer

OpenStudy (anonymous):

so the third?

OpenStudy (neo92):

if timo86m is correct. third one z da answer.

OpenStudy (anonymous):

do u thinlk im right?

OpenStudy (amistre64):

not paralell, but same slope and intercepts

OpenStudy (anonymous):

so its the third poone right

OpenStudy (neo92):

im not sure abt tht parellel thing. wait Il ask sm 1

OpenStudy (amistre64):

spose one equation is a scaled version of the other, they are then equal by all accounts

OpenStudy (amistre64):

ax+by = c ::: n(ax + by = c)

OpenStudy (anonymous):

haha so the third one? cause thats what i keep getting

OpenStudy (neo92):

yep. da third one z da nly possible answer.

OpenStudy (amistre64):

lets see (7y = x + 3) *3 -3x + 21y = 9 21y = 3x + 9 -3x + 21y = 9 i like it